# Frame-Wise Breath Detection with Self-Training: An Exploration of Enhancing Breath Naturalness in Text-to-Speech

### v1

@@ Abstract
Developing Text-to-Speech (TTS) systems that can synthesize natural breath is essential for human-like voice agents but requires extensive manual annotation of breath positions in training data. To this end, we propose a self-training method for training a breath detection model that can automatically detect breath positions in speech. Our method trains the model using a large speech corpus and involves: 1) annotation of limited breath sounds utilizing a rule-based approach, and 2) iterative augmentation of these annotations through pseudo-labeling based on the model's predictions. Our detection model employs Conformer blocks with down-/up-sampling layers, enabling accurate frame-wise breath detection. We investigate its effectiveness in multi-speaker TTS using text transcripts with detected breath marks. The results indicate that using our proposed model for breath detection and breath mark insertion synthesizes breath-contained speech more naturally than a baseline model.
@@ 1 Introduction
To enhance the naturalness of breath in synthetic speech, a widely adopted strategy is to detect breath sounds in speech data and insert corresponding breath marks into the text data for TTS training [2, 4]. Then the breath marks are predicted and inserted during the inference phase [3]. However, conventional breath detection methods come with several limitations. They generally fall into two categories: one relies on acoustic feature analysis for rule-based detection [11, 2], and the other manually annotates the breath data and trains machine learning models for detection [12, 13]. The former, though not requiring manual data annotation, tends to compromise on accuracy; while the latter, despite achieving higher accuracy, fails to yield a universal and robust detection model without extensive annotation across large speech corpora. To our knowledge, these methods have only been applied in small speech corpora.
In this study, we primarily concentrate on breath detection to overcome existing limitations. We introduce an innovative training approach that eliminates the need for manual annotation of breath in the training dataset. Our approach is a hybrid one that utilizes the advantages of previous rule-based and machine-learning-based approaches with a self-training method. Moreover, we propose a powerful frame-wise breath detection model, trained on a large multi-speaker speech corpus. Our method starts with a comprehensive acoustic feature analysis for breath extraction. We not only verify the effectiveness of several features from previous works but also put forward additional features. Based on these features, we develop a rule-based approach to automatically extract and annotate a sufficient number of breath sounds within the training set, aiming to train the breath detection model. Our proposed breath detection model incorporates Conformer [14] and bidirectional long short-term memory (BiLSTM). It is optimized with downsampling and upsampling modules, enabling frame-wise detection while reducing computational costs. Through a self-training methodology [15, 16], the model is trained efficiently with both annotated and unannotated data, leading to superior performance. Experimental results demonstrate that our proposed model significantly outperforms the baseline model in breath detection. When training the TTS model with speech and text transcripts augmented with detected breath marks from the proposed model, the synthetic breath sounds exhibit enhanced naturalness. Speech samples are available online [^1] https://ydqmkkx.github.io/breath-detection/.
@@
  • Our proposed model achieves frame-wise breath detection with finer time resolution compared to the baseline model's frame-level detection.
@@ 2.1 Data, acoustic features, and annotation
Montreal Forced Aligner (MFA) [18] is utilized for aligning text transcripts with speech and detecting pauses. These pauses include several kinds of PINTs, mainly silence, breath, and tongue clicks. To evaluate our breath detection methods, we manually annotated the breath sounds within these pauses. The annotation was conducted on randomly selected sentences from the ``dev-clean'' and ``test-clean'' subsets of LibriTTS-R, thereby forming respective validation and test sets of breath detection experiments, the statistics of which are detailed in Table 1.
@@ 2.3 Self-training method
The self-training process, as outlined in Algorithm 1, begins with the training of the initial detector $D_{\theta}^{0}$. In subsequent iterations, the breath and non-breath sets are augmented with pseudo-labels ($\hat{B}$ and $\hat{U}$). The pseudo-labels are generated based on the detector's predictions and dynamic thresholds ($\alpha^{k}$ and $\beta^{k}$). Specifically, the frames with a predictive probability above $\alpha^{k}$ are assigned to $\hat{B}$, while those with a predictive probability below $\beta^{k}$ are included in $\hat{U}$. These pseudo-labels are used to progressively refine the detector's training. The iterative training continues until the detector’s performance declines on the validation set, yielding the best-performing detector $D_{\theta}^{k-1}$.
@@ 3.1 Breath detection experiments
Configurations: The ``train-clean-100'' and ``train-other-500'' subsets of LibriTTS-R corpus were used as the training set. The breath and non-breath sets were extracted and annotated following the rule-based breath detection outlined in Section 2.1. Speech signals were sampled at a rate of 16,000 Hz and processed to extract log-scaled Mel-spectrograms with 128 Mel bands, to serve as input features. The training was conducted on one NVIDIA A100 GPU with the AdamW optimizer [23]. The linear learning rate scheduler was utilized, with the initial 10% training steps as the warm-up period. The learning rate increased from 0 to the peak during the warm-up period and then decreased to 0 linearly. During each training iteration, both baseline and proposed models were trained for 10 epochs with a batch size of 64 and a peak learning rate of $2\times 10^{-5}$.
@@
Results and analysis: For this detection task, Intersection over Union (IoU) was adopted as the primary evaluation metric, complemented by precision and recall for more detailed analysis. The optimal predictive threshold was determined at the point where the model maximized IoU for breath detection in the validation set. This threshold was then applied to calculate these metrics in the test set. The experimental results are listed in Table 3, which reveal the following findings:
  • From Table 3(a), the proposed model consistently outperformed the baseline model, demonstrating its robustness. In addition, both of the two models achieved their peak IoU after the 3rd iteration of self-training, where the models were considered as the best-performing ones. Then the training finished after the 4th iteration, where both models improved the precision further but degraded IoU and recall a little.
  • From Table 3(b), the absence of either ZCR or VMS in the input, especially ZCR, significantly reduced the model's performance. Additionally, incorporating the non-breath set into the training was also proved to be critical. Furthermore, instead of improving, the proposed model's performance declined after the 1st iteration without pseudo-labeling, highlighting the significance of the self-training method.
@@ 3.2 TTS experiments
Configurations: To effectively train the TTS models with sufficient pauses, the speech utterances in the ``train-clean-360'' subset of LibriTTS-R were segmented into clips ranging from 5 to 10 seconds for training and evaluation. We employed VITS [24] as the backbone TTS model and utilized the baseline and proposed model (after the 3rd iteration in Table 3) for breath detection. Based on the detection results, the breath marks were inserted into the text transcripts of both training and test sets, resulting in three TTS models: VITS, VITS w/ baseline, and VITS w/ proposed. These models were trained on two NVIDIA A100 GPUs for 200 epochs, with the same configurations of optimizer and learning rate scheduler as detailed in Section 3.1. The batch size was 160 and the peak learning rate was $5\times 10^{-4}$. During inference, the two key hyper-parameters of VITS, ``noise_scale_w'' and ``noise_scale'' were set to 0.0 and 0.667, respectively.
@@ 4 Discussion
Through multi-speaker training with breath detection and breath mark insertion, we observed that VITS w/ proposed could synthesize breath sounds to some extent for speakers who lacked them in the training data. Specifically, in the ``train-clean-360'' subset, speaker ``3630'' and ``1811'' lacked breath sounds. VITS w/ proposed could synthesize the breath sounds for both speakers, whereas VITS w/ baseline only managed this for ``1811''. We speculate that this was because the baseline model incorrectly detected some silence as breath sounds in the training data of speaker ``3630'', leading the TTS model to misinterpret silence as its breath sounds.

### v2

@@ Abstract
Developing Text-to-Speech (TTS) systems that can synthesize natural breath is essential for human-like voice agents but requires extensive manual annotation of breath positions in training data. To this end, we propose a self-training method for training a breath detection model that can automatically detect breath positions in speech. Our method trains the model using a large speech corpus and involves: 1) annotation of limited breath sounds utilizing a rule-based approach, and 2) iterative augmentation of these annotations through pseudo-labeling based on the model’s predictions. Our detection model employs Conformer blocks with down-/up-sampling layers, enabling accurate frame-wise breath detection. We investigate its effectiveness in multi-speaker TTS using text transcripts with detected breath marks. The results indicate that using our proposed model for breath detection and breath mark insertion synthesizes breath-contained speech more naturally than a baseline model.
@@ 1 Introduction
To enhance the naturalness of breath in synthetic speech, a widely adopted strategy is to detect breath sounds in speech data and insert corresponding breath marks into the text data for TTS training [2, 4]. Then, the breath marks are predicted and inserted during the inference phase [3]. However, conventional breath detection methods come with several limitations. They generally fall into two categories: one relies on acoustic feature analysis for rule-based detection [11, 2], and the other manually annotates the breath data and trains machine learning models for detection [12, 13]. The former, though not requiring manual data annotation, tends to compromise on accuracy; while the latter, despite achieving higher accuracy, fails to yield a universal and robust detection model without extensive annotation across large speech corpora. To our knowledge, these methods have only been applied in small speech corpora.
In this study, we primarily concentrate on breath detection to overcome existing limitations. We introduce an innovative training approach that eliminates the need for manual annotation of breath in the training dataset. Our approach is a hybrid one that utilizes the advantages of previous rule-based and machine-learning-based approaches with a self-training method. Moreover, we propose a powerful frame-wise breath detection model trained on a large multi-speaker speech corpus. Our method starts with a comprehensive acoustic feature analysis for breath extraction. We not only verify the effectiveness of several features from previous works but also put forward additional features. Based on these features, we develop a rule-based approach to automatically extract and annotate a sufficient number of breath sounds within the training set, aiming to train breath detection models. Our proposed breath detection model incorporates Conformer [14] and bidirectional long short-term memory (BiLSTM). It is optimized with downsampling and upsampling modules, enabling frame-wise detection while reducing computational costs. Through a self-training methodology [15, 16], the model is trained efficiently with both annotated and unannotated data, leading to superior performance. Experimental results demonstrate that our proposed model significantly outperforms the baseline model in breath detection. When training the TTS model with speech and text transcripts augmented with detected breath marks from the proposed model, the synthetic breath sounds exhibit enhanced naturalness. Speech samples and code are available [^1] https://ydqmkkx.github.io/breath-detection/.
@@
  • Our proposed model achieves frame-wise breath detection with finer time resolution compared to the baseline model’s frame-level detection.
@@ 2.1 Data, acoustic features, and annotation
Montreal Forced Aligner (MFA) [18] is utilized for aligning text transcripts with speech and detecting pauses. These pauses include several kinds of PINTs, mainly silence, breath, and tongue clicks. To evaluate our breath detection methods, we manually annotated the breath sounds within these pauses. The annotation was conducted on randomly selected sentences from the “dev-clean” and “test-clean” subsets of LibriTTS-R, thereby forming respective validation and test sets of breath detection experiments, the statistics of which are detailed in Table 1.
@@ 2.3 Self-training method
The self-training process, as outlined in Algorithm 1, begins with the training of the initial detector $D_{\theta}^{0}$. In subsequent iterations, the breath and non-breath sets are augmented with pseudo-labels ($\hat{B}$ and $\hat{U}$). The pseudo-labels are generated based on the detector’s predictions and dynamic thresholds ($\alpha^{k}$ and $\beta^{k}$). Specifically, the frames with a predictive probability above $\alpha^{k}$ are assigned to $\hat{B}$, while those with a predictive probability below $\beta^{k}$ are included in $\hat{U}$. These pseudo-labels are used to progressively refine the detector’s training. The iterative training continues until the detector’s performance declines on the validation set, yielding the best-performing detector $D_{\theta}^{k-1}$.
@@ 3.1 Breath detection experiments
Configurations: The “train-clean-100” and “train-other-500” subsets of LibriTTS-R corpus were used as the training set. The breath and non-breath sets were extracted and annotated following the rule-based breath detection outlined in Section 2.1. Speech signals were sampled at a rate of 16,000 Hz and processed to extract log-scaled Mel-spectrograms with 128 Mel bands, to serve as input features. The training was conducted on one NVIDIA A100 GPU with the AdamW optimizer [23]. The linear learning rate scheduler was utilized, with the initial 10% training steps as the warm-up period. The learning rate increased from 0 to the peak during the warm-up period and then decreased to 0 linearly. During each training iteration, both baseline and proposed models were trained for 10 epochs with a batch size of 64 and a peak learning rate of $2\times 10^{-5}$.
@@
Results and analysis: For this detection task, Intersection over Union (IoU) was adopted as the primary evaluation metric, complemented by precision and recall for more detailed analysis. The optimal predictive threshold was determined at the point where the model maximized IoU for breath detection in the validation set. This threshold was then applied to calculate these metrics in the test set. The experimental results are listed in Table 3(b), which reveal the following findings:
  • From Table 3(b)(a), the proposed model consistently outperformed the baseline model, demonstrating its robustness. In addition, both of the two models achieved their peak IoU after the 3rd iteration of self-training, where the models were considered as the best-performing ones. Then the training finished after the 4th iteration, where both models improved the precision further but degraded IoU and recall a little.
  • From Table 3(b)(b), the absence of either ZCR or VMS in the input, especially ZCR, significantly reduced the model’s performance. Additionally, incorporating the non-breath set into the training was also proved to be critical. Furthermore, instead of improving, the proposed model’s performance declined after the 1st iteration without pseudo-labeling, highlighting the significance of the self-training method.
@@ 3.2 TTS experiments
Configurations: To effectively train the TTS models with sufficient pauses, the speech utterances in the “train-clean-360” subset of LibriTTS-R were segmented into clips ranging from 5 to 10 seconds for training and evaluation. We employed VITS [24] as the backbone TTS model and utilized the baseline and proposed model (after the 3rd iteration in Table 3(b)) for breath detection. Based on the detection results, the breath marks were inserted into the text transcripts of both training and test sets, resulting in three TTS models: VITS, VITS w/ baseline, and VITS w/ proposed. These models were trained on two NVIDIA A100 GPUs for 200 epochs, with the same configurations of optimizer and learning rate scheduler as detailed in Section 3.1. The batch size was 160 and the peak learning rate was $5\times 10^{-4}$. During inference, the two key hyper-parameters of VITS, “noise_scale_w” and “noise_scale” were set to 0.0 and 0.667, respectively.
@@ 4 Discussion
Through multi-speaker training with breath detection and breath mark insertion, we observed that VITS w/ proposed could synthesize breath sounds to some extent for speakers who lacked them in the training data. Specifically, in the “train-clean-360” subset, speaker “3630” and “1811” lacked breath sounds. VITS w/ proposed could synthesize the breath sounds for both speakers, whereas VITS w/ baseline only managed this for “1811”. We speculate that this was because the baseline model incorrectly detected some silence as breath sounds in the training data of speaker “3630”, leading the TTS model to misinterpret silence as its breath sounds.